OpenStudy (anonymous):

how do i write answer Find the equation of the circle with radius 6 and center at the origin. x^2+y^2 or (x+y)^2

OpenStudy (anonymous):

\[x^2 +y^2 = 36\]

OpenStudy (anonymous):

Thanks mate

OpenStudy (anonymous):

\[x^2 + y^2 = r^2 \]

OpenStudy (anonymous):

You're welcome